3. (9 marks) Evaluate the integral / V9 – 22 - dx
Answer #1

| 59-22 dx = 39 - 82 1 / 2 de Integration by parts Suvda = u fudor - Jl de svd x) dx - 0 Here (u= V9 - 82), V = 1 1 1 2 de =(vozidel doc = ufudre - flat Suda) dx substituting all values in = 1965) - S +) dox to ☺ 9-12 usir F10 by parts 9n v gare? Vg